# Tilus: A Virtual Machine for Arbitrary Low-Precision GPGPU Computation in LLM Serving

### v2

@@ Abstract.
Serving Large Language Models (LLMs) is critical for AI-powered applications but demands substantial computational resources, particularly in memory bandwidth and computational throughput. Low-precision computation has emerged as a key technique to improve efficiency while reducing resource consumption. Existing approaches for generating low-precision kernels are limited to weight bit widths that are powers of two and suffer from suboptimal performance due to high-level GPU programming abstractions. These abstractions restrict critical optimizations, such as fine-grained register management and optimized memory access patterns, which are essential for efficient low-precision computations. In this paper, we introduce a virtual machine (VM) designed for General-Purpose GPU (GPGPU) computing, enabling support for low-precision data types with arbitrary bit widths while maintaining GPU programmability. The proposed VM features a thread-block-level programming model, a hierarchical memory space, a novel algebraic layout system, and extensive support for diverse low-precision data types. VM programs are compiled into highly efficient GPU programs with automatic vectorization and instruction selection. Extensive experiments demonstrate that our VM efficiently supports a full spectrum of low-precision data types, and outperforms state-of-the-art low-precision kernels on their supported types. Compared to existing compilers like Triton and Ladder, as well as hand-optimized kernels such as QuantLLM and Marlin, our VM achieves performance improvements of 1.75×1.75$\times$1.75 ×, 2.61×2.61$\times$2.61 ×, 1.29×1.29$\times$1.29 × and 1.03×1.03$\times$1.03 ×, respectively.
@@ 1. Introduction
Quantization (Frantar et al., 2022; Lin et al., 2024; Ashkboos et al., 2024; Chee et al., 2023; Tseng et al., 2024; Liu et al., 2025) has emerged as a leading method for enhancing the efficiency of LLM serving. By reducing the bit-width of model parameters and activations, quantization reduces weight storage, DRAM bandwidth usage, and achieves faster computation. For instance, A16W4 quantization (16-bit activation and 4-bit weight) reduces the DRAM consumption and throughput by 4×4$\times$4 × compared with A16W16 scheme, thus reducing the time to generate a token by about 4×4$\times$4 × (Frantar et al., 2024). However, the state-of-the-art 4-bit quantization methods (Ashkboos et al., 2024; Chee et al., 2023; Liu et al., 2025) still suffer from non-negligible accuracy degradation. While using 5- to 7-bit quantization could mitigate this accuracy loss (Agrawal et al., 2024; Xia et al., 2024), the lack of efficient GPU kernels for these bit widths has hindered their adoption. Generating optimized kernels for hardware-unfriendly bit widths (e.g., 3, 5, 6, and 7) remains an open problem.
@@
To address these limitations, compiler-based approaches (Chen et al., 2018; Wang et al., 2024; Tillet et al., 2019) have been proposed to automate kernel generation. Among them, Triton (Tillet et al., 2019) simplifies the GPGPU programming through a tile-based model. A Triton program defines the computation of a thread block over tensor tiles. However, Triton lacks built-in support for low-precision data types, requiring users to implement low-level bitwise operations manually. Additionally, it does not expose the full GPU memory hierarchy, limiting optimization opportunities for low-precision LLM inference. Ladder (Wang et al., 2024), on the other hand, extends TVM’s scheduling system (Chen et al., 2018) to support low-precision computation but is restricted to data types with bit widths that are powers of two. Moreover, its primitives cannot express crucial optimizations such as software pipelining (Huang et al., 2023), leading to suboptimal performance, particularly for batch sizes greater than one during LLM decoding with continuous batching (Yu et al., 2022) enabled.
To address these challenges, we propose Tilus, a GPGPU virtual machine with dedicated support for low-precision computation. Tilus abstracts GPU program execution into thread-block-level instructions, simplifying GPGPU programming while exposing hierarchical memory spaces for fine-grained manipulation of sub-tensors in on-chip memory. This dual approach enables efficient handling of arbitrary-precision data types while reducing the complexity of GPU programming. To achieve these goals, Tilus introduces: (1) an algebraic layout system that specifies how tensor elements within a tile are distributed across GPU threads. This layout system enables flexible reinterpretation of low-precision tiles in registers as tiles with hardware-friendly data types for efficient processing; (2) a thread-block-level programming model with fine-grained memory management, providing explicit control over data movement, placement, and computation across different levels of the GPU memory hierarchy; and comprehensive support for (3) arbitrary low-precision data types including signed integers, unsigned integers, and floating-point numbers with 1 to 8 bits.
Extensive experiments show that Tilus extends the spectrum of efficient low-precision kernels to support arbitrary bit widths (1-8) and data type kinds (e.g., integer and floating numbers). At the same time, Tilus outperforms the state-of-the-art compiler Triton (Tillet et al., 2019), Ladder (Wang et al., 2024), and hand-crafted kernels from QuantLLM (Xia et al., 2024) and Marlin (Frantar et al., 2024) by 1.75×1.75$\times$1.75 ×, 2.61×2.61$\times$2.61 ×, 1.29×1.29$\times$1.29 × and 1.03×1.03$\times$1.03 × on their supported low-precision kernels, respectively.
@@
  • We propose a GPGPU virtual machine with dedicated support for low-precision computation, addressing the critical coverage (e.g., support 5 to 7 bits quantization) and performance gap in existing approaches.
  • Inside the virtual machine, we introduce a novel layout system, a thread-block-level programming model with hierarchical memory space, and support low-precision data types with arbitrary bit widths.
  • Through extensive evaluation, we demonstrate that Tilus generates a full spectrum of highly efficient low-precision kernels, outperforming state-of-the-art solutions across their supported kernels by up to 2.6×2.6$\times$2.6 ×.
@@ 2.1. LLM Serving and Quantization
In large language model (LLM) serving, inference consists of two stages: prefill and decode. The prefill stage processes the input prompt to establish context, while the decode stage iteratively generates output tokens based on prior tokens. Among all layers of LLMs, matrix multiplications dominate computation time and memory consumption, making their optimization crucial for efficient LLM serving. Quantization (Dettmers et al., 2024; Frantar et al., 2022) improves their efficiency by reducing model weights and activations to lower-precision formats, such as 8-bit or 4-bit integers. It reduces memory usage, bandwidth requirements, and inference latency while trying to maintain model accuracy. While 4-bit quantization provides significant computational savings, state-of-the-art methods (Ashkboos et al., 2024; Chee et al., 2023; Liu et al., 2025) still suffer from accuracy degradation. Increasing precision to 5-bit, 6-bit, or 7-bit quantization (Agrawal et al., 2024; Xia et al., 2024) can help preserve accuracy while maintaining efficiency, but these bit widths lack optimized GPU support, limiting their adoption. Current GPU architectures and software stacks primarily optimize for power-of-two bit widths (e.g., 4-bit and 8-bit), making arbitrary bit widths computationally inefficient. However, demand for flexible quantization is growing, as 4-bit can be too aggressive for some models while 8-bit wastes resources. Supporting a broader spectrum of bit widths enables better accuracy-efficiency trade-offs in LLM serving, driving the need for new kernel generation techniques that can efficiently handle non-standard low-precision formats (e.g., those with 3, 5, 6, 7 bit widths) on modern GPUs.
2.2. The GPGPU Programming
General-Purpose GPU (GPGPU) programming enables parallel computation by organizing tasks within a structured execution and memory hierarchy (Nickolls and Dally, 2010). The execution hierarchy begins with the thread, the smallest unit of execution, which performs instructions independently with its own registers and local memory. Threads are grouped into thread blocks, which enable data sharing through shared memory and support synchronized execution. A grid consists of multiple independent thread blocks, enabling large-scale parallelism by organizing thousands or millions of threads. The GPU memory hierarchy comprises registers, shared memory, and global memory. Registers provide the fastest and thread-private storage. Shared memory is accessible by all threads within a thread block and faster than global memory. Global memory is accessible across the entire grid with high latency. This structure allows for highly efficient parallel execution by leveraging both the execution and memory hierarchies.
@@ 2.3. The GPGPU Languages and Compilers
2.3.1. GPGPU Virtual Machine and Languages
GPGPU programming involves various languages and compilers that balance hardware abstraction with control. Low-level languages like SASS (NVIDIA Corporation, 2024c) and CDNA3 (AMD Corporation, 2024a) offer direct hardware access for fine-grained optimizations but require deep architectural knowledge. Slightly higher in abstraction, NVIDIA’s PTX (NVIDIA Corporation, 2024b) serves as an intermediate representation, linking high-level languages such as CUDA (NVIDIA Corporation, 2024a) to GPU-specific instructions while preserving optimization flexibility. High-level languages like CUDA (NVIDIA Corporation, 2024a) and HIP (AMD Corporation, 2024b) simplify programming by extending C programming language. Despite these languages, GPGPU programming remains complex. It is constrained by hardware-specific memory and computation hierarchies and requires workload-specific optimizations. To address these challenges, researchers have introduced higher-level languages and compilers, classified into two categories: procedure-oriented compilers, which simplify programming through abstractions beyond CUDA (NVIDIA Corporation, 2024a), and schedule-oriented compilers, which optimize computation-hardware mappings via declarative scheduling primitives.
2.3.2. Procedure-Oriented Compilers
This type of compilers (Lattner et al., 2021; Hagedorn et al., 2023; Ding et al., 2023; Tillet et al., 2019) enables programmers to write kernels directly, offering abstractions to simplify the process. Triton (Tillet et al., 2019), for instance, introduces the tile programming model, where thread block behavior is defined programmatically, and tiles replace scalars as the basic data type. This approach combines programming simplicity with high-performance kernel generation, making Triton widely adopted. However, Triton lacks native support for low precision data types like uint4. Handling these types requires manually unpacking sub-byte data from larger storage types (e.g., uint32(Hoque et al., 2024). Additionally, Triton does not expose the GPU memory hierarchy, limiting programmers’ control over data loading and memory scope usage, which complicates performance optimization for low-precision kernels. These limitations result in inefficient low-precision kernel execution. Figure 1(a) illustrates the inefficiencies in Triton-generated low-precision kernels, using a uint4 weight loading pipeline as an example. The process includes four steps: 1 weights are asynchronously copied from global memory to shared memory using pipelined cp.async instructions (Huang et al., 2023); 2 shared memory data is loaded into registers; 3 unpacking and casting operations are performed; and 4 the register tensor layout is converted to meet the requirements of tensor core instructions. Among these, step 4 is a major bottleneck due to the reliance on shared memory for layout conversion, which incurs significant overhead.
@@ 2.3.3. Schedule-Oriented Compilers
Schedule-oriented compilers decouple computation from scheduling to optimize the computation-to-hardware mappings. Halide (Ragan-Kelley et al., 2013) pioneered this approach, which was later extended by TVM (Chen et al., 2018) and subsequent works (Vaswani et al., 2017; Zheng et al., 2020a, 2022a; Shao et al., 2024; Huang et al., 2023; Feng et al., 2023; Wang et al., 2024) in the domain of deep learning. Among them, Ladder (Wang et al., 2024) is the first one to support low-precision computation by introducing dedicated primitives to pack low-precision data (e.g., 4-bit integers) into larger types (e.g., 8-bit integers). However, Ladder (Wang et al., 2024) has two limitations. First, it cannot handle non-power-of-two bit widths efficiently due to type-level packing, packing low-precision types into storage types. Second, its primitive-style scheduling prevents optimizations like software pipelining (Huang et al., 2023), resulting in suboptimal performance. Figure 1 (b) illustrates the weight loading process in Ladder’s low-precision kernels. This process includes 1 loading weights from global memory to registers without pipelining; 2 vectorized casting; 3 storing the cast results in shared memory; and finally 4 using the ldmatrix instruction to load weights from shared memory to registers for subsequent tensor core operations. The lack of pipelining between weight loading and computation significantly hinders performance.
@@
Figure 1. The weight loading pipeline of Triton, Ladder, and our approach. The tensors could be in either global memory (GMEM), shared memory (SMEM), and registers (REGS).
Figure 2. This figure provides a concrete example of how the GPGPU virtual machine is used to implement low-precision matrix multiplication (FP16 $\times$ INT6). Figure (a) illustrates the virtual machine program, highlighting key features such as the algebraic layout system (Section 4), thread-block-level instructions (Section 5), and efficient low-precision data handling. Figure (b) illustrates the kernel’s data flow, emphasizing tensor movement across the memory hierarchy and intermediate operations such as tensor reinterpretation and type casting. A similar weight-loading strategy can be applied to arbitrary type widths (Section 6). Finally, Figure (c) demonstrates register tensor reinterpretation, showing how tensors with compatible bit distributions across threads (e.g., 24 bits per thread) can be efficiently reinterpreted into different data types and layouts.
@@ 3.1. Key Ideas
Our work introduces a novel GPGPU virtual machine (VM) specifically designed to overcome the challenges of programming efficient low-precision deep learning kernels. The VM natively supports low-precision data types with arbitrary bit widths ranging from 1 to 8, enabling efficient weight loading and computation. Figure 1 (c) shows the VM’s weight loading pipeline, using uint4 as an example. It begins with 1 a pipelined asynchronous memory copy from global memory to shared memory, followed by 2 loading the register tensor from shared memory. Next, it 3 reinterprets the register tensor into a different data type and layout at no cost, and then 4 performs vectorized casting. This pipeline achieves superior efficiency compared to the other methods in Figure 1, as it eliminates layout conversion (unlike Triton (Tillet et al., 2019)) and incorporates pipelining (unlike Ladder (Wang et al., 2024)). More importantly, our pipeline is versatile, making our work the first to seamlessly support arbitrary low-precision data types with bit widths ranging from 1 bit to 8 bits.
@@
To achieve this efficiency, our design is built on several key ideas. A GPGPU Virtual Machine for Flexible Optimization: our decision to implement a GPGPU VM stems from the need for greater flexibility in GPU programming. Unlike standard loop-based transformations, our VM allows programmers to directly implement and fine-tune optimizations that go beyond traditional loop transformations. This flexibility is critical for low-precision computations, where fine-grained control over execution strategies can lead to significant performance gains. A Thread-Block-Level Programming Model with Hierarchical Memory Spaces: our VM explicitly exposes the GPU memory hierarchy, including registers, shared memory, and global memory, that are abstracted away in existing solutions like Triton (Tillet et al., 2019). By granting programmers fine-grained control over data placement and movement, our approach enables memory pipelining and eliminates unnecessary layout conversions, as shown in Figure 1. An Algebraic Layout System: we introduce an algebraic layout system that precisely defines how elements within a register tensor are distributed among threads. This structured representation simplifies the construction, analysis, and interpretation of tensor layouts. Notably, it enables seamless reinterpretation of low-precision register tensors into standard data types, as demonstrated in Step 3 of Figure 1(c). Native Support for Arbitrary Low-Precision Data Types: our VM provides built-in support for a wide range of low-precision data types, including both signed and unsigned integers and floating-point numbers with bit widths from 1 to 8. Supported types include int2 to int8, uint1 to uint8, and float3 to float8, with arbitrary exponent and mantissa distribution for floating-point types. These innovations collectively enhance the programmability, efficiency, and flexibility of low-precision kernel development on modern GPUs. We chose not to extend Triton (Tillet et al., 2019) because its programming model inherently abstracts away tensor layouts, making it incompatible with our approach of explicit layout control. Similarly, Ladder (Wang et al., 2024) relies on type-level packing, whereas Tilus employs tile-level reinterpretation, making the two fundamentally incompatible. The next section presents an example of low-precision matrix multiplication within our virtual machine.
3.2. An Example of Virtual Machine Program

### v3

# Tilus: A Tile-Level GPGPU Programming Language for Low-Precision Computation

@@ Abstract.
Serving Large Language Models (LLMs) is critical for AI-powered applications, yet it demands substantial computational resources, particularly in memory bandwidth and computational throughput. Low-precision computation has emerged as a key technique to improve efficiency while reducing resource consumption. Existing approaches for generating low-precision kernels are limited to weight bit widths that are powers of two and suffer from suboptimal performance because of high-level GPU programming abstractions. These abstractions restrict critical optimizations, such as fine-grained register management and optimized memory access patterns, that are essential for efficient low-precision computations. In this paper, we introduce Tilus, a domain-specific language designed for General-Purpose GPU (GPGPU) computing that supports low-precision data types with arbitrary bit widths from 1 to 8 while maintaining GPU programmability. Tilus features a thread-block-level programming model, a hierarchical memory space, a novel algebraic layout system, and extensive support for diverse low-precision data types. Tilus programs are compiled into highly efficient GPU programs through automatic vectorization and instruction selection. Extensive experiments demonstrate that Tilus efficiently supports a full spectrum of low-precision data types, and outperforms state-of-the-art low-precision kernels. Compared to existing compilers such as Triton and Ladder, as well as hand-optimized kernels such as QuantLLM and Marlin, Tilus achieves performance improvements of: $1.75\times$, $2.61\times$, $1.29\times$ and $1.03\times$, respectively. We open-source Tilus at https://github.com/NVIDIA/tilus.
GPU; programming language; parallel computation; low-precision computation; quantization
$c$s: Computing methodologies Parallel programming languagesccs: Computing methodologies Machine learningjournalyear: 2026copyright: rightsretainedconference: xxx; xxx; xxxbooktitle: xxx
@@ 1. Introduction
Quantization (Frantar et al., 2022; Lin et al., 2024; Ashkboos et al., 2024; Chee et al., 2023; Tseng et al., 2024; Liu et al., 2025) has emerged as a leading method for enhancing the efficiency of LLM serving. By reducing the bit-width of model parameters and activations, quantization reduces weight storage, DRAM bandwidth usage, and achieves faster computation. For instance, A16W4 quantization (16-bit activation and 4-bit weight) reduces DRAM consumption and throughput by $4\times$ compared to the A16W16 scheme, thereby reducing the time to generate a token by about $4\times$ (Frantar et al., 2024). However, the state-of-the-art 4-bit quantization methods (Ashkboos et al., 2024; Chee et al., 2023; Liu et al., 2025) still suffer from non-negligible accuracy degradation. While using 5- to 7-bit quantization can mitigate this accuracy loss (Agrawal et al., 2024; Xia et al., 2024), the lack of efficient GPU kernels for these bit widths has hindered their adoption. Generating optimized kernels for hardware-unfriendly bit widths (e.g., 3, 5, 6, and 7) remains an open problem.
@@
To address these limitations, compiler-based approaches (Chen et al., 2018; Wang et al., 2024; Tillet et al., 2019) have been proposed to automate kernel generation. Among them, Triton (Tillet et al., 2019) simplifies the GPGPU programming through a tile-based model. A Triton program defines the computation of a thread block for tensor tiles. However, Triton lacks built-in support for low-precision data types, and thus requires users to implement low-level bitwise operations manually. Additionally, it does not expose the full GPU memory hierarchy, limiting optimization opportunities for low-precision LLM inference. Ladder (Wang et al., 2024), on the other hand, extends TVM’s scheduling system (Chen et al., 2018) to support low-precision computation but is restricted to data types with bit widths that are powers of two. Moreover, its primitives cannot express crucial optimizations such as software pipelining (Huang et al., 2023), leading to suboptimal performance, particularly for batch sizes greater than one during LLM decoding with continuous batching (Yu et al., 2022) enabled.
In response to these challenges, we propose Tilus, a tile-level GPGPU programming language backed by a virtual machine with dedicated support for low-precision computation. Tilus abstracts GPU program execution into thread-block-level instructions, simplifying GPGPU programming while exposing hierarchical memory spaces for fine-grained manipulation of sub-tensors in on-chip memory. This dual approach enables efficient handling of arbitrary-precision data types while reducing the complexity of GPU programming. To achieve these goals, Tilus introduces: (1) an algebraic layout system that specifies how tensor elements within a tile are distributed across GPU threads. This layout system provides a unified way to represent the storage of tile elements among different threads, making it possible to expose the registers to the kernel developers and simplify the code generation during compilation. (2) a thread-block-level programming model with fine-grained memory management, providing explicit control over data movement, placement, and computation across different levels of the GPU memory hierarchy; and comprehensive support for (3) arbitrary low-precision data types including signed integers, unsigned integers, and floating-point numbers with bit widths ranging from 1 to 8.
Extensive experiments show that Tilus extends the spectrum of efficient low-precision kernels to support arbitrary bit widths (from 1 to 8) and data type kinds (e.g., integer and floating-point numbers). At the same time, Tilus outperforms the state-of-the-art compilers, such as Triton (Tillet et al., 2019) and Ladder (Wang et al., 2024), as well as hand-crafted kernels from QuantLLM (Xia et al., 2024) and Marlin (Frantar et al., 2024), achieving performance improvements of: $1.75\times$, $2.61\times$, $1.29\times$ and $1.03\times$, respectively.
@@
  • We propose a GPGPU programming language Tilus with dedicated support for low-precision computation, addressing the critical bit-width coverage and performance gap in existing approaches.
  • Within Tilus, we introduce a novel layout system, a thread-block-level programming model with hierarchical memory space, and support low-precision data types with arbitrary bit-width from 1 to 8.
  • Through extensive evaluation, we demonstrate that Tilus generates a full spectrum of highly efficient low-precision kernels, achieving up to $2.6\times$ performance improvement over state-of-the-art approaches on their supported kernels.
@@ 2.1. LLM Serving and Quantization
LLM serving consists of two inference stages: prefill and decode. The prefill stage processes the input prompt to establish context, while the decode stage iteratively generates output tokens based on prior tokens. Key LLM layers include multi-head attention, feed-forward networks, and layer normalization (Vaswani et al., 2017a). Among these, matrix multiplications dominate computation time and memory consumption, making their optimization crucial for efficient LLM serving. Quantization (Dettmers et al., 2024; Frantar et al., 2022) improves their efficiency by reducing model weights and activations to lower-precision formats, such as 8-bit or 4-bit integers. It reduces memory usage, bandwidth requirements, and inference latency while aiming to preserve model accuracy. While 4-bit quantization provides significant computational savings, state-of-the-art methods (Ashkboos et al., 2024; Chee et al., 2023; Liu et al., 2025) still suffer from accuracy degradation. Increasing precision to 5-bit, 6-bit, or 7-bit quantization (Agrawal et al., 2024; Xia et al., 2024) can help preserve accuracy while maintaining efficiency, but these bit widths lack optimized GPU support, limiting their adoption. Current GPU architectures and software stacks primarily optimize for power-of-two bit widths (e.g., 4-bit and 8-bit), making arbitrary bit widths computationally inefficient. However, demand for flexible quantization is growing, as 4-bit can be too aggressive for some models while 8-bit wastes resources. Supporting a broader spectrum of bit widths enables better accuracy-efficiency trade-offs in LLM serving, driving the need for new kernel generation techniques that can efficiently handle non-standard low-precision formats (e.g., those with 3, 5, 6, 7 bit widths) on modern GPUs.
2.2. GPGPU Programming
General-Purpose GPU (GPGPU) programming enables parallel computation by organizing tasks within a structured execution and memory hierarchy (Nickolls and Dally, 2010). The execution hierarchy begins with the thread, the smallest unit of execution, which performs instructions independently, using its own registers and local memory. Threads are grouped into thread blocks, which enable data sharing through shared memory and support synchronized execution. A grid consists of multiple independent thread blocks, enabling large-scale parallelism by organizing thousands or millions of threads. The GPU memory hierarchy comprises registers, shared memory, and global memory. Registers provide the fastest and thread-private storage. Shared memory is accessible by all threads within a thread block and faster than global memory. Global memory is accessible across the entire grid with high latency. This structure allows for highly efficient parallel execution by leveraging both the execution and memory hierarchies.
@@ 2.3. The GPGPU Languages and Compilers
2.3.1. GPGPU Programming Languages
GPGPU programming involves various languages and compilers that balance hardware abstraction with control. Low-level languages like SASS (NVIDIA Corporation, 2024c) and CDNA3 (AMD Corporation, 2024a) offer direct hardware access for fine-grained optimizations but require deep architectural knowledge. Slightly higher in abstraction, NVIDIA’s PTX (NVIDIA Corporation, 2024b) serves as an intermediate representation that links high-level languages like CUDA (NVIDIA Corporation, 2024a) to GPU-specific instructions while preserving optimization flexibility. High-level languages like CUDA (NVIDIA Corporation, 2024a) and HIP (AMD Corporation, 2024b) simplify programming by extending the C programming language. Despite these languages, GPGPU programming remains complex. It is constrained by hardware-specific memory and computation hierarchies and requires workload-specific optimizations. To address these challenges, researchers have introduced higher-level languages and compilers, classified into two categories: tile-oriented compilers, which simplify programming through abstractions beyond CUDA (NVIDIA Corporation, 2024a), and schedule-oriented compilers, which optimize computation-hardware mappings via declarative scheduling primitives.
2.3.2. Tile-Oriented Compilers
This type of compilers, such as Graphene (Hagedorn et al., 2023), Hidet (Ding et al., 2023), and Triton (Tillet et al., 2019), enables programmers to write kernels directly, offering abstractions like tile types or tile-level task/element distribution to simplify the process. Triton (Tillet et al., 2019), for instance, introduces the tile programming model, where thread block behavior is defined programmatically, and tiles replace scalars as the basic data type. This approach combines programming simplicity with high-performance kernel generation, making Triton widely adopted. However, Triton lacks native support for low precision data types like uint4. Handling these types requires manually unpacking sub-byte data from larger storage types (e.g., uint32(Hoque et al., 2024). Additionally, Triton does not expose the GPU memory hierarchy, limiting programmers’ control over data loading and memory scope usage, which complicates performance optimization for low-precision kernels. These limitations result in inefficient low-precision kernel execution. Figure 1(a) illustrates the inefficiencies in Triton-generated low-precision kernels, using a uint4 weight loading pipeline as an example. The process includes four steps: 1 weights are asynchronously copied from global memory to shared memory using pipelined cp.async instructions (Huang et al., 2023); 2 shared memory data is loaded into registers; 3 unpacking and casting operations are performed; and 4 the register tensor layout is converted to meet the requirements of tensor core instructions. Among these, step 4 is a major bottleneck due to the reliance on shared memory for layout conversion, which incurs significant overhead.
@@ 2.3.3. Schedule-Oriented Compilers
Schedule-oriented compilers decouple computation from scheduling to optimize the computation-to-hardware mappings. Halide (Ragan-Kelley et al., 2013) pioneered this approach, which was later extended by TVM (Chen et al., 2018) and subsequent works (Vaswani et al., 2017b; Zheng et al., 2020a, 2022a; Shao et al., 2024; Huang et al., 2023; Feng et al., 2023; Wang et al., 2024) in the domain of deep learning. Among them, Ladder (Wang et al., 2024) is the first one to support low-precision computation by introducing dedicated primitives to pack low-precision data (e.g., 4-bit integers) into larger types (e.g., 8-bit integers). However, Ladder (Wang et al., 2024) has two limitations. First, it cannot handle non-power-of-two bit widths efficiently due to type-level packing, packing low-precision types into storage types. Second, its primitive-style scheduling prevents optimizations like software pipelining (Huang et al., 2023), resulting in suboptimal performance. Figure 1 (b) illustrates the weight loading process in Ladder’s low-precision kernels. This process includes 1 loading weights from global memory to registers without pipelining; 2 vectorized casting; 3 storing the cast results in shared memory; and finally 4 using the ldmatrix instruction to load weights from shared memory to registers for subsequent tensor core operations. This lack of pipelining between weight loading and computation significantly hinders performance.
@@
Figure 1. The weight loading pipeline of Triton, Ladder, and our approach. The tensors could be in global memory (GMEM), shared memory (SMEM), or registers (REGS).
Figure 2. This figure provides a concrete example of how the Tilus is used to implement low-precision matrix multiplication (FP16 $\times$ INT6). Figure (a) illustrates the virtual machine program, highlighting key features such as the algebraic layout system (Section 4), thread-block-level instructions (Section 6), and efficient low-precision data support. Figure (b) illustrates the kernel’s data flow, emphasizing tensor movement across the memory hierarchy and intermediate operations such as tensor reinterpretation and type casting. This similar weight-loading strategy can be applied to arbitrary type widths (Section 7). Finally, Figure (c) demonstrates register tensor reinterpretation, showing how tensors with compatible bit distributions across threads (e.g., 24 bits per thread) can be efficiently reinterpreted into different data types and layouts.
@@ 3.1. Key Ideas
Our work introduces a domain-specific language, Tilus, that provides fine-grained control over shared memory and registers, making it possible to program efficient low-precision deep learning kernels. Tilus supports low-precision data types with arbitrary bit widths ranging from 1 to 8, enabling efficient weight loading and computation. Figure 1 (c) shows the weight loading pipeline of a Tilus program, using uint4 as an example. It begins with 1 a pipelined asynchronous memory copy from global memory to shared memory, followed by 2 loading the register tensor from shared memory. Next, it 3 reinterprets the register tensor into a different data type and layout at no cost, before finally 4 performing vectorized casting. This pipeline achieves superior efficiency compared to the other methods in Figure 1, as it eliminates layout conversion (unlike Triton (Tillet et al., 2019)) and incorporates pipelining (unlike Ladder (Wang et al., 2024)). More importantly, our pipeline is generic, making our work the first to seamlessly support arbitrary low-precision data types with bit widths ranging from 1 to 8 bits.
@@
To achieve this efficiency, our design is built on several key ideas. A GPGPU Virtual Machine: every Tilus program is a program for an abstract GPGPU virtual machine (VM) that contains an instruction set. This decision stems from the need for greater flexibility in GPU programming. By abstracting GPU functionalities, such as memory loading and computation, into instructions, it becomes easier to add support for new architectural features while keeping support for older ones. A Thread-Block-Level Programming Model with Hierarchical Memory Spaces: The underlying VM explicitly exposes the GPU memory hierarchy — including registers, shared memory, and global memory — that existing solutions like Triton (Tillet et al., 2019) abstract away. By granting programmers fine-grained control over data placement and movement, our approach enables memory pipelining and eliminates unnecessary layout conversions, as shown in Figure 1. An Algebraic Layout System: we introduce an algebraic layout system that precisely defines how elements within a register tensor are distributed among threads. This structured representation simplifies the construction, analysis, and interpretation of tensor layouts. Notably, it enables seamless reinterpretation of low-precision register tensors into standard data types, as demonstrated in Step 3 of Figure 1(c). Native Support for Arbitrary Low-Precision Data Types: Tilus provides built-in support for a wide range of low-precision data types, including both signed and unsigned integers and floating-point numbers with bit widths from 1 to 8. Supported types include int2 to int8, uint1 to uint8, and float3 to float8, with arbitrary exponent and mantissa distribution for floating-point types. These innovations collectively enhance the programmability, efficiency, and flexibility of low-precision kernel development on modern GPUs. We chose not to extend Triton (Tillet et al., 2019) because its programming model inherently abstracts away tensor layouts, making it incompatible with our approach of explicit layout control. Similarly, Ladder (Wang et al., 2024) relies on type-level packing, whereas Tilus employs tile-level reinterpretation, making the two fundamentally incompatible. The next section presents a Tilus-programmed example of low-precision matrix multiplication.
3.2. An Example of Tilus Program